# Quantum latent distributions in deep generative models

### v1

@@ Appendix A Complexity Proofs
We define, with $\mathcal{C}$, the space of sampling problems of the form $\mathcal{S}=(P_{z_{n}})_{n\in\mathbb{N}}$ (where each $P_{z_{n}}$ is a probability distribution over $\mathbb{R}^{n}$) such that we can sample from distributions $(\hat{P}_{z_{n}})_{n\in\mathbb{N}}$ with $W(P_{z_{n}},\hat{P}_{z_{n}})<\epsilon$ [^1] WW denotes the Wasserstein metric using a classical algorithm in $\text{Poly}(n,1/\epsilon)$ time.
@@
We define, with $\mathcal{Q}$, the space of sampling problems of the form $\mathcal{S}=(P_{z_{n}})_{n\in\mathbb{N}}$ (where each $P_{z_{n}}$ is a probability distribution over $\mathbb{N}^{n}$) such that we can sample from a distribution $\hat{P}_{g_{n}(z_{n})}$ such that $\|P_{z_{n}}-\hat{P}_{g_{n}(z_{n})}\|<\epsilon$ using a quantum circuit in $\text{Poly}(n,1/\epsilon)$ time, and there exists no classical algorithm that can do the same.
@@
Generative Model Setting: We define our collection of latent distributions as a collection $(P_{z_{n}})_{n\in\mathbb{N}}$, where $P_{z_{n}}$ is a distribution over $\mathbb{R}^{n}$. We will consider collections of neural network architectures $(g_{n})_{n\in\mathbb{N}}$ where $g_{n}$ takes as input a vector of length $n$. In this section, we will assume these architectures are all efficiently classically implementable in the sense that there exists a $\text{Poly}(n)$ classical algorithm that applies the collection of functions. We will generally use $(P_{g_{n}(z_{n})})_{n\in\mathbb{N}}$ to denote the pushforward distributions, where:
@@
Proof: Let $(\hat{P}_{z_{n}})_{n\in\mathbb{N}}$ be the collection of efficiently classically sample-able approximations guaranteed to exist by $(P_{z_{n}})_{n\in\mathbb{N}}\in\mathcal{C}$. Let $\hat{P}_{g_{n}(z_{n})}$ denote the distribution sampled from by applying $g_{n}$ to values sampled according to $\hat{P}_{z_{n}}$. We note:
@@
Theorem 1: Let the neural networks $(g_{n})_{n\in\mathbb{N}}$ be invertible such that there is a classical algorithm that applies $g_{n}^{-1}$ in $\text{Poly}(n)$ time and further, $g_{n}^{-1}$ is Lipschitz continuous with constant $c$. Let $(P_{z_{n}})_{n\in\mathbb{N}}$ be in $\mathcal{Q}$. Then, sampling from the collection of pushforward distributions $(P_{g_{n}(z_{n})})_{n\in\mathbb{N}}$ is not in $\mathcal{C}$.
Proof: For a contradiction, we assume that the sampling task defined by $(P_{g_{n}(z_{n})})_{n\in\mathbb{N}}$ is in $\mathcal{C}$. Let $(\hat{P}_{g_{n}(z_{n})})_{n\in\mathbb{N}}$ denote the classically efficiently sample-able approximation guaranteed to exist by the definition of $\mathcal{C}$. Let $\hat{P}_{z_{n}}$ denote the distribution sampled by applying $g_{n}^{-1}$ to values sampled according to $\hat{P}_{g_{n}(z_{n})}$. We claim that $W(P_{z_{n}},\hat{P}_{z_{n}})<c\epsilon$. Indeed,
@@
Corollary 1: Let us be in the same setting as described in Theorem 1. Let $(\hat{P}_{z_{n}})_{n\in\mathbb{N}}\in\mathcal{C}$ be a collection of latent distributions we can sample from exactly in Poly($n$) time. Let $(\hat{P}_{g_{n,\epsilon}(z_{n})})_{n\in\mathbb{N},\epsilon\in\mathbb{R}}$ be the resulting collection of pushforward distributions.
There exists $\epsilon\in\mathbb{R}$, $n\in\mathbb{N}$ such that:
@@
Proof: By assumption, we can sample from $(\hat{P}_{g_{n,\epsilon}(z_{n})})$ in Poly($n,1/\epsilon)$ time. If $\forall n,\epsilon$, we have $W(\hat{P}_{g_{n,\epsilon}(z_{n})},P_{g_{n}(z_{n})})<\epsilon$, then $(P_{g_{n}(z_{n})})_{n\in\mathbb{N}}\in\mathcal{C}$. This contradicts Theorem 1. Thus $\exists\,n,\epsilon$ such that :
$W(\hat{P}_{g_{n}(z_{n}))},P_{g_{n}(z_{n})})\geq\epsilon$
Precisely, the result demonstrated is that, no matter the choice of classical latent distribution and efficiently implementable neural network architecture, it is not possible to simulate the performance of a specific set of GAN architectures utilising latent distributions generated by quantum circuits.
@@ B.1 Boson sampling theory
where Perm indicates the matrix permanent and $U_{k,T}$ is the submatrix of $U$ obtained by taking the first $k$ columns of $U$ and by repeating each row $i$ $t_{i}$ times (if $t_{i}=0$ then row $i$ is not in $U_{k,T}$). The permanent of an $n$ by $n$ matrix $M$ with elements $x_{i,j}$ is defined as
$\text{Perm}(M)=\sum_{\sigma\in S_{n}}\prod_{i=i}^{n}x_{i,\sigma_{i}}$
where $S_{n}$ is the set of all permutations of the numbers 1 to $n$. Although this expression is similar to that of a matrix determinant, unlike a determinant calculating the permanent of a matrix is in general a computationally hard problem. Indeed, if $M$ is an arbitrary real or complex matrix then this problem is #P-hard [11]. Since calculating a permanent is hard, [11] show that the problem of sampling from the output distribution of a boson sampler, even approximately, is also hard. The classical complexity of simulating a boson sampler increases exponentially with the number of photons, such that current classical computers cannot simulate more than a few tens of photons [45].
B.2 Current state of the art
Over the last few years, improvements in photon source, circuit and detector technology have led to several boson sampling experiments at a scale that is a challenge for classical computers. For instance, [29] performed a boson sampling experiment in which up to 76 photons were measured in 100 channels using a fixed interference circuit. Later, [14] performed a boson sampling experiment in an optical fiber loop system involving up to 219 photons in 216 channels. However, classical algorithms were later discovered that can leverage the high loss levels in these systems to simulate them [59]. More recently, a demonstration of boson sampling with thousands of photons was performed, with the authors claiming that this cannot be simulated with any existing classical algorithm [63].
B.3 Boson samplers with delay lines
@@
Boson samplers based on optical delay lines have been proposed and demonstrated as an experimentally feasible route to large-scale boson sampling [14, 51, 52]. In these systems, illustrated in figure 5, a single photon source is used to produce a sequence of identical single photons, which then interfere with other in a sequence of optical delay lines (or "loops"). The resulting superposition state is measured by a single photon number resolving detector. The main advantage of this approach is that only a single photon source and detector are required, even though many photons and time-bins may be involved. Moreover, only three loops of different lengths are required to reach the classically hard to simulate regime [51, 52]. Furthermore, in [14] and in the ORCA Computing PT-2, the coupling parameter between each loop and each input/output is re-programmable between each time step using fast electro-optic modulators. This opens the door to the possibility of jointly training these parameters with the parameters of a neural network.
B.4 ORCA Computing PT-2 processor
The boson sampling system used in our experiments is an ORCA Computing PT-2, which is a commercially available loop-based boson sampling system. It consists of a photon source based on parametric downconversion, two sequential optical delay lines of the same length, and photon-number resolving superconducting nanowire detectors. It can support up to 16 photons interfering in 32 time-bins. The coupling parameters between each optical delay line and the input/output channels are fully programmable.
@@ B.4 ORCA Computing PT-2 processor
To run our experiments, we collected a dataset of 500k samples with fixed set of randomly initialized coupling parameters, which took 40 minutes on this system. The system we used is hosted at the UK National Quantum Computing Centre.
@@ Appendix C Experiments on 2D mixtures of Gaussians
These experiments used a WGAN-GP framework, where both the generator and discriminator are feedforward neural networks using two hidden layers with 256 neurons and LeakyReLU activations. All latent distributions were size-16. Both the distinguishable and indistinguishable samples were produced by randomly initialized simulated photonic systems with a 3-loop architecture in a 1-3-9 configuration. The models were trained for 5000 training steps, and all distributions were centered to have a mean value of 0 before being injected into the generator. Though figure 2 presents results for only a single seed, the experiments were repeated for 5 seeds with repeatable behavior. When training for 10,000 iterations, we observed that the latent distribution with distinguishable photons was often able to close the gap with the indistinguishable photons, which implies that the generator is able to overcome the difference between those distributions on this dataset with further training. However, further training did not improve results with Gaussian and Bernoulli latent distributions.
Appendix D Experiments on a toy dataset
@@ Appendix E QM9 experiments
Our implementation started from the repository at https://github.com/kfzyqin/Implementation-MolGAN-PyTorch. We made a few changes to lead to improved performance. First, the MolGAN generator was modified by feeding affine transforms of the latent code z to all layers of the feedforward network, similar to the approach taken in [65]. The activation functions used in the generator were also changed to LeakyReLU. Moreover, instead of using 3 hidden layers, we used 5 hidden layers with sizes 64, 176, 288, 400 and 512.
@@
Figure 6: FCD scores during training for different latents of $z_{dim}$=16. The mean and standard error of the mean are calculated from 20 random seeds.
@@ E.1 Note on circuit randomization
In general, to ensure that performance differences arise from general properties of the latent space distributions and not from specific circuit realizations, the optical circuits for both distinguishable and indistinguishable photons are re-sampled for every seed. This is the case for all experiments with a toy dataset, and also for all size-16 and most size-32 latents on the QM9 dataset.
@@ Appendix F DDGANs
The idea behind Denoising Diffusion GANS ([55]) is that to have a faster reverse process we need to go back multiple steps at a time, in which case the assumption that the denoising distributions are Gaussian fails. A solution is to perform the reverse process using more complex multimodal distributions. To achieve this, a full multimodal conditional GAN is used at each step of the denoising diffusion process. The generative process is performed using 4 denoising steps ($T=4$). For a given (fixed) Gaussian noise input $\textbf{x}_{T}$, different images $\textbf{x}_{0}$ are obtained using different samples from the latent distribution $z$. The combination of the initial Gaussian noise and samples from the latent determine the final result.
The boson and distinguishable sampler latents for these experiments were produced by two unstructured optical circuits, with transfer matrices described by Haar-random unitary matrices, with 24 photons in 48 channels.
We use the code in https://github.com/NVlabs/denoising-diffusion-gan to run these experiments. The code was minimally modified to use the boson and distinguishable samplers as latents for the generator during training and inference. We used the commands in the repo to train and evaluate the models with the only difference being that the latent space dimension in the original model was $z_{dim}=256$. Some example outputs are shown in figure 8.
@@ Appendix G Flow matching
We build on the Optimal-Transport variant of Conditional Flow Matching from [27], as implemented in https://github.com/atong01/conditional-flow-matching. To reconcile the dimensionality mismatch between the quantum latent distribution and our two-dimensional data, we introduce an untrained decoder that maps quantum latents to 2D source samples. Specifically, we first apply an affine transformation, scaling and shifting each latent vector by its global mean and variance, and then project the result through a Kaiming-normal–initialized kernel. This ensures that the resultant source distribution remains comparable in scale to the standard Gaussian source distribution. The decoder’s parameters remain fixed throughout training, and all other training procedures follow the original codebase. During sampling, we draw from the quantum latent distribution, decode each sample, and then perform the usual flow-matching routine.
@@ Appendix H Negative results
In this section, we present additional results for experiments in which we found that the performance of a model was not improved by using a quantum latent distribution.
@@ H.2 QM9 experiments
We also found that small changes to the model and training regime used in our QM9 experiments could make a significant difference and erase a large part of the observed performance differences. Using the same model but with a higher learning rate of $10^{-3}$ (instead of $10^{-4}$), a permutation probability of 0.2, and 10,000 training steps (instead of 20,000) led to the results shown in table 8 for size-16 latents. We see that the results are worse overall, and that there is also a much smaller difference between the different latents. This indicates that performance differences achieved using a different latent distributions are sensitive to model and training parameters dataset.
@@
Latent Type $z_{dim}$ FCD $\downarrow$ # Valid & unique $\uparrow$ # Novel $\uparrow$

### v2

@@ Appendix A Complexity Proofs
We define, with $\mathcal{C}$, the space of sampling problems of the form $\mathcal{S}=(P_{z_{n}})_{n\in\mathbb{N}}$ (where each $P_{z_{n}}$ is a probability distribution over $\mathbb{R}^{n}$) such that we can sample from distributions $(\hat{P}_{z_{n}})_{n\in\mathbb{N}}$ with $W(P_{z_{n}},\hat{P}_{z_{n}})<\epsilon$ using a classical algorithm in $\text{Poly}(n,1/\epsilon)$ time. Here, $W$ denotes the Wasserstein metric
@@
We define, with $\mathcal{Q}$, the space of sampling problems of the form $\mathcal{S}=(P_{z_{n}})_{n\in\mathbb{N}}$ (where each $P_{z_{n}}$ is a probability distribution over $\mathbb{N}^{n}$) such that we can sample from a distribution $\hat{P}_{z_{n}}$ with $\|P_{z_{n}}-\hat{P}_{z_{n}}\|<\epsilon$ using a quantum circuit in $\text{Poly}(n,1/\epsilon)$ time, and there exists no classical algorithm that can do the same. Here, $\|\cdot\|$ denotes total variation distance.
@@
Generative Model Setting: We denote our collection of latent distributions as $(P_{z_{n}})_{n\in\mathbb{N}}$, where $P_{z_{n}}$ is a distribution over $\mathbb{R}^{n}$. We will consider collections of neural network architectures $(g_{n})_{n\in\mathbb{N}}$ where $g_{n}$ takes as input a vector of length $n$. In this section, we will assume these architectures are all efficiently classically implementable in the sense that there exists a $\text{Poly}(n)$ classical algorithm that applies the collection of functions. We will generally use $(P_{g_{n}(z_{n})})_{n\in\mathbb{N}}$ to denote the pushforward distributions, where:
@@
Proof: Let $(\hat{P}_{z_{n}})_{n\in\mathbb{N}}$ be the collection of efficiently classically sampleable approximations guaranteed to exist by $(P_{z_{n}})_{n\in\mathbb{N}}\in\mathcal{C}$. Let $\hat{P}_{g_{n}(z_{n})}$ denote the distribution sampled from by applying $g_{n}$ to values sampled according to $\hat{P}_{z_{n}}$. We note:
@@
Theorem 1: Let the neural networks $(g_{n})_{n\in\mathbb{N}}$ be invertible such that there is a classical algorithm that applies $g_{n}^{-1}$ in $\text{Poly}(n)$ time and further, each $g_{n}^{-1}$ is Lipschitz continuous with constant $c$. Let $(P_{z_{n}})_{n\in\mathbb{N}}$ be in $\mathcal{Q}$. Then, sampling from the collection of pushforward distributions $(P_{g_{n}(z_{n})})_{n\in\mathbb{N}}$ is not in $\mathcal{C}$.
Proof: For a contradiction, we assume that the sampling task defined by $(P_{g_{n}(z_{n})})_{n\in\mathbb{N}}$ is in $\mathcal{C}$. Let $(\hat{P}_{g_{n}(z_{n})})_{n\in\mathbb{N}}$ denote the classically efficiently sampleable approximation guaranteed to exist by the definition of $\mathcal{C}$. Let $\hat{P}_{z_{n}}$ denote the distribution sampled by applying $g_{n}^{-1}$ to values sampled according to $\hat{P}_{g_{n}(z_{n})}$. We claim that $W(P_{z_{n}},\hat{P}_{z_{n}})<c\epsilon$. Indeed,
@@
Corollary 1: Let us be in the same setting as described in Theorem 1. Let $(\hat{P}_{z_{n}})_{n\in\mathbb{N}}\in\mathcal{C}$ be a collection of latent distributions we can sample from exactly in Poly($n$) time. Let $(\hat{P}_{g_{n,\epsilon}(z_{n})})_{n\in\mathbb{N},\epsilon\in\mathbb{R}}$ be the resulting collection of pushforward distributions. There exists $\epsilon\in\mathbb{R}$, $n\in\mathbb{N}$ such that:
@@
Proof: To work towards a contradiction, we assume that $D^{G_{n,\epsilon}}(\hat{P}_{z_{n}},P_{g_{n}(z_{n})})<\epsilon$ holds $\forall n,\epsilon$. We then pick the functions $g_{n,\epsilon}\in G_{n,\epsilon}$ such that $W(\hat{P}_{g_{n,\epsilon}(z_{n})},P_{g_{n}(z_{n})})<\epsilon$. As we can sample from $(\hat{P}_{g_{n,\epsilon}(z_{n})})$ in Poly($n,1/\epsilon)$ time, $(P_{g_{n}(z_{n})})_{n\in\mathbb{N}}\in\mathcal{C}$. This contradicts Theorem 1.
Precisely, the result demonstrated is that, no matter the choice of latent distribution and efficiently implementable neural network architecture, it is not possible to efficiently classically simulate the performance of a specific set of GAN architectures whose latent distributions are generated by quantum circuits.
@@ B.1 Boson sampling theory
where Perm indicates the matrix permanent and $U_{k,T}$ is the submatrix of $U$ obtained by taking the first $k$ columns of $U$ and repeating each row $i$ $t_{i}$ times (if $t_{i}=0$ then row $i$ is not in $U_{k,T}$). The permanent of an $n$ by $n$ matrix $M$ with elements $x_{i,j}$ is defined as
$\text{Perm}(M)=\sum_{\sigma\in S_{n}}\prod_{i=1}^{n}x_{i,\sigma_{i}}$
where $S_{n}$ is the set of all permutations of the numbers 1 to $n$. Although this expression is similar to that of a matrix determinant, unlike a determinant calculating the permanent of a matrix is in general a computationally hard problem. Indeed, if $M$ is an arbitrary real or complex matrix then this problem is #P-hard. Since calculating a permanent is hard, [1] show that the problem of sampling from the output distribution of a boson sampler, even approximately, is also hard. The classical complexity of simulating a boson sampler increases exponentially with the number of photons, such that current classical computers cannot simulate more than a few tens of photons [13].
B.2 Boson samplers with delay lines
@@
Boson samplers based on optical delay lines have been proposed and demonstrated as an experimentally feasible route to large-scale boson sampling [39, 16, 45]. In these systems, illustrated in figure 3, a single photon source is used to produce a sequence of identical single photons, which then interfere with each other in a sequence of optical delay lines (or "loops"). The resulting superposition state is measured by a single photon number resolving detector. The main advantage of this approach is that only a single photon source and detector are required, even though many photons and time-bins may be involved. Moreover, only three loops of different lengths are required to reach the classically hard to simulate regime [16, 45]. Furthermore, in [39] and in the ORCA Computing PT-2, the coupling parameter between each loop and each input/output is re-programmable between each time step using fast electro-optic modulators. This opens the door to the possibility of jointly training these parameters with the parameters of a neural network.
B.3 ORCA Computing PT-2 processor
The boson sampling system used in our experiments in section IV.2.3 is an ORCA Computing PT-2, which is a commercially available loop-based boson sampling system. It consists of a photon source based on parametric downconversion, two sequential optical delay lines of the same length, and photon-number resolving superconducting nanowire detectors. It supports up to 16 photons interfering in 32 time-bins. The coupling parameters between each optical delay line and the input/output channels are fully programmable.
@@ B.3 ORCA Computing PT-2 processor
To run our experiments, we collected a dataset of 500k samples with a fixed set of randomly initialized coupling parameters. Collecting this data took 40 minutes on this system. The system we used is hosted at the UK National Quantum Computing Centre.
@@ Appendix C Experiments on 2D mixtures of Gaussians
These experiments used a WGAN-GP framework, where both the generator and discriminator are feedforward neural networks using two hidden layers with 256 neurons and LeakyReLU activations. All latent distributions were size-16. Both the distinguishable and indistinguishable samples were produced by randomly initialized simulated photonic systems with a 3-loop architecture in a 1-3-9 configuration. The models were trained for 5000 training steps, and all distributions were centered to have a mean value of 0 before being injected into the generator. Though figure 1 presents results for only a single seed, the experiments were repeated for 5 seeds with qualitatively repeatable behavior. When training for 10,000 iterations, we observed that the latent distribution with distinguishable photons was often able to close the gap with the indistinguishable photons, which implies that the generator is able to overcome the difference between those distributions on this dataset with further training. However, further training did not improve results with Gaussian and Bernoulli latent distributions. This is in line with the results from [38] on a similar toy dataset, where it was found that surprisingly deep neural networks were required to generate this dataset from an initial Gaussian distribution.
Appendix D Experiments on a synthetic quantum dataset
@@ Appendix E QM9 experiments
Our implementation started from the repository at https://github.com/kfzyqin/Implementation-MolGAN-PyTorch. We made a few changes to lead to improved performance. First, the MolGAN generator was modified by feeding affine transforms of the latent code z to all layers of the feedforward network, similar to the approach taken in [29]. The activation functions used in the generator were also changed to LeakyReLU. Moreover, instead of using 3 layers, we used 5 layers with sizes 64, 176, 288, 400 and 512.
@@
Figure 4: FCD scores during training for different latents of $d_{z}$=16. The mean and standard error of the mean are calculated from 20 random seeds.
@@ E.1 Note on circuit randomization
In general, to ensure that performance differences arise from general properties of the latent space distributions and not from specific circuit realizations, the optical circuits for both distinguishable and indistinguishable photons are re-sampled for every seed. This is the case for all experiments with the synthetic quantum dataset, and also for all size-16 and most size-32 latents on the QM9 dataset.
@@ Appendix F DDGANs
Latent diffusion models generally make Gaussianity assumptions on the latent distribution [50], such that the use of a quantum distribution is not straightforward. However, it has been observed that the reverse process, when considered over large time steps, involves modeling a complex non-Gaussian distribution, for which a GAN can be well suited [64], therefore we focus on Denoising Diffusion GANs (DDGANs), which use the standard forward diffusion process but use a GAN to perform the reverse process. They achieve results comparable to standard diffusion models with faster inference speeds, with a model that is easier to train than other GANs. This work aligns with the hypothesis [25] that at least part of the observed improvement in diffusion models over GANs may stem from the additional computation at inference due to the recursive sampling process rather than a fundamental paradigm difference.
In DDGANs, a full multimodal conditional GAN is used at each step of the denoising diffusion process. The generative process is performed using 4 denoising steps ($T=4$). For a given (fixed) Gaussian noise input $\textbf{x}_{T}$, different images $\textbf{x}_{0}$ are obtained using different samples from the latent distribution $z$. The combination of the initial Gaussian noise and samples from the latent determine the final result.
We investigated unconditional generation on the CIFAR-10 dataset [31] for different latent spaces, where the original model used a $d_{z}=256$ latent space, our experiments used a $d_{z}=48$ latent space due to the classical complexity of simulating larger quantum circuits. The quantum and distinguishable sampler latents for these experiments were produced by two unstructured optical circuits, with transfer matrices described by Haar-random unitary matrices, with 24 photons in 48 channels.
We used the code in https://github.com/NVlabs/denoising-diffusion-gan to run these experiments. The code was minimally modified to use the boson and distinguishable samplers as latents for the generator during training and inference. We used the commands in the repo to train and evaluate the models. Some example outputs are shown in figure 6. Though these experiments do not exhibit a performance advantage with a quantum latent distribution, they demonstrate the feasibility of this approach and motivate further exploration.
@@ Appendix G Flow matching
To illustrate the feasibility of a quantum latent distribution as a source distribution, Figure 7 represents a small-scale model trained using the method from [55] to map 2D-projected quantum latent vectors to the "moons" dataset. We build on the Optimal-Transport variant of Conditional Flow Matching from [55], as implemented in https://github.com/atong01/conditional-flow-matching. To reconcile the dimensionality mismatch between the quantum latent distribution and our two-dimensional data, we introduce an untrained decoder that maps quantum latents to 2D source samples. Specifically, we first apply an affine transformation, scaling and shifting each latent vector by its global mean and variance, and then project the result through a Kaiming-normal–initialized kernel. This ensures that the resultant source distribution remains comparable in scale to the standard Gaussian source distribution. The decoder’s parameters remain fixed throughout training, and all other training procedures follow the original codebase. During sampling, we draw from the quantum latent distribution, decode each sample, and then perform the usual flow-matching routine.
@@ Appendix H Negative results
In this section, we present additional results for experiments in which we found that the performance of a GAN model was not improved by using a quantum latent distribution.
@@ H.2 QM9 experiments
We also found that small changes to the model and training regime used in our QM9 experiments could make a significant difference and erase a large part of the observed performance differences. Using the same model but with a higher learning rate of $10^{-3}$ (instead of $10^{-4}$), a permutation probability of 0.2, and 10,000 training steps (instead of 20,000) led to the results shown in table 8 for size-32 latents. There is no statistically significant difference between the latent distributions in this regime. The Gaussian distribution exhibits improved performance compared to the results from table 2, though still significantly worse than the best results using the size-16 latents. These results indicate that performance differences achieved using a different latent distribution are sensitive to the model and to the training parameters.
@@
Latent Type $d_{z}$ FCD $\downarrow$ # Valid & unique $\uparrow$ # Novel $\uparrow$